OpenStudy (anonymous):

Write the equation in slope intercept form m=-2 (6,8)

OpenStudy (ddcamp):

Start by putting this in the point slope form: y-y₁ = m(x-x₁) Then, solve for y by adding y₁ to both sides, leaving you with: y = mx - mx₁ + y₁

OpenStudy (anonymous):

What's the answer